OpenStudy (anonymous):

Find the leading coefficient and the degree of the following polynomial: p(x)=-3(x-1)^4(3x+1)(5-2x)^3 The lelading term would be (x-1) or would it be 3(x-1)? then wouldn't the leading constant be -3? or 1?

OpenStudy (anonymous):

then wouldn't my degree be 4?

OpenStudy (jdoe0001):

hmm I think you have to expand the whole thing first

OpenStudy (anonymous):

expand?

OpenStudy (ranga):

\[p(x) = -3(x-1)^{4} (3x + 1) (5 - 2x)^{3}\] Is this how the expression appears?

OpenStudy (anonymous):

yep

OpenStudy (jdoe0001):

well... just the higher exponent binomial, you may need to expand only

OpenStudy (ranga):

Look at the leading term in each sub-expression and just multiply them. That would suffice.

OpenStudy (jdoe0001):

though using the binomial theorem, is rather visible what his the coefficient

OpenStudy (ranga):

(-3) * (x^4) * (3x) * (-2x)^3 should be sufficient to get the leading term of the product.

OpenStudy (anonymous):

I do that for every polynomial to get the leading term? How would I then get the degree?

OpenStudy (jdoe0001):

notice what ranga is doing, after expanding each binomial, he's just picking the highest exponent term and multiplying, which will yield the highest exponent from their product

OpenStudy (jdoe0001):

in short, you had to expand all 3 binomials, then multiply them, and grab the highest term for the degree of the polynomial

OpenStudy (jdoe0001):

but a short way to do it, is like ranga did, noticing the binomial theorem, just pick the highest term from each binomial, and multiply them

OpenStudy (ranga):

(-3) * (x^4) * (3x) * (-2x)^3 = (-3)(3)(-2)^3 * (x^4)(x)(x)^3 = (-3)(3)(-8) * (x^4)(x)(x^3) = 72 * x^8 is the leading term in the product.

OpenStudy (anonymous):

okay thank you I am going to look over this for a min

OpenStudy (anonymous):

whats the binomial therom?

OpenStudy (ranga):

Sure. The long way to do it will be to multiply the whole thing out using the binomial theorem to expand the fourth power and the third power, then gather all the like terms together and rearrange the polynomial from the highest degree to the lowest. But that will be too consuming. Since all they are interested in is just the leading term you can use the above shortcut. Binomial theorem lets you find: (a + b)^n. If n = 2, it is easy: a^2 + 2ab + b^2 But as n increases, such as n = 3, 4, 5, 6, etc. it is hard to remember all those formula. Instead one can remember the binomial theorem and do it for any n.

OpenStudy (anonymous):

is that attachment a different leading coefficient?

OpenStudy (ranga):

No.\[\Large a _{n}x^{n}\]is the leading term and \[\Large a_{n}\]is the leading coefficient. But in the attachment the polynomial has been nicely arranged from the highest degree to the lowest degree and so the leading term is easily found. But here we will have to multiply out everything and gather the like terms,simplify, rearrange from the highest to the lowest degree and only then we will see what the polynomial is. So we have to do this multiplication: -3 * (x-1) * (x-1) * (x-1) * (x-1) * (3x+1) * (5-2x) * (5-2x) * (5-2x). But that is too time consuming. We can use binomial theorem but still it will be too time consuming. So I showed you a shortcut if all they want is just the leading term.

OpenStudy (anonymous):

ohhh okay. Thank you im going to learn it right now.

OpenStudy (anonymous):

what happens to the 1 in (3x+1) and the 5 in (5-2x)^3?

OpenStudy (ranga):

If you really multiply it all out, this is what the polynomial will look like: 72x^8 - 804x^7 + 3666x^6 - 8739x^5 + 11481x^4 - 7746x^3 + 1620x^2 + 825x - 375 As you can see the leading coefficient is 72x^8.

OpenStudy (anonymous):

oh my goodness .....

OpenStudy (ranga):

So instead of doing the multiplication we looked at each sub-expression of -3(x-1)^4(3x+1)(5-2x)^3: (-3) * (x-1)^4 * (3x+1) * (5-2x)^3 The constant (-3) will remain as it is. (x-1)^4 will have x^4 as the leading term (3x+1) will have 3x as the leading term (5-2x)^3 will have (-2x)^3 as the leading term So if we multiply all of the leading terms of each sub-expression we will know what the leading term of the entire polynomial will be.

OpenStudy (anonymous):

how did you get -3?

OpenStudy (ranga):

-3 is in the problem.It is the first thing after the = sign: p(x)=-3(x-1)^4(3x+1)(5-2x)^3

OpenStudy (anonymous):

why would (-2x)^3 be the leading term and not 5?

OpenStudy (ranga):

The phrase "leading term" may be misleading. They use that phrase because normally polynomials are supposed to be arranged from the highest degree to the lowest degree. If that is done, then the phrase "leading term" will be synonymous with "the first term". But sometimes polynomials are not arranged properly and so the leading term will not be the first term but we have to find what is the term that has the highest degree. In (5-2x)^3 inside the parenthesis the polynomial is not arranged properly. It should have been (-2x + 5)^3 and then you can easily see that -2x is the leading term. The leading term is always the term that has the highest degree.

OpenStudy (anonymous):

sorry one more questions how does -2x have the highest degree?

OpenStudy (ranga):

the polynomial -2x + 5 has two terms. The constant is like degree zero. x^1 is degree 1. So -2x is the leading term.

OpenStudy (anonymous):

ohh okay thank you = )

OpenStudy (ranga):

sure. hope things are a bit clearer now. If you want another problem for practice: Find the leading term of: 3 * (1 + 2x)^2 * (2 - 5x)

OpenStudy (anonymous):

okay brb

OpenStudy (ranga):

ok

OpenStudy (anonymous):

the leading term would be 3,2x,and 5x ?

OpenStudy (ranga):

3,2x,and -5x (don't forget the sign)

OpenStudy (anonymous):

yes -5x. =)

OpenStudy (ranga):

And you have to multiply all the leading terms remembering the middle term is squared.

OpenStudy (anonymous):

2x^2?

OpenStudy (ranga):

The middle term will be (2x)^2

OpenStudy (ranga):

The leading term of 3 * (1 + 2x)^2 * (2 - 5x) will be: 3 * (2x)^2 * (-5x) (all you have to is just drop everything but the leading term in each sub-expression.)

OpenStudy (anonymous):

okay and I take that multiply it together and get the leading coefficient

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

nice thank you

OpenStudy (ranga):

you are welcome.